# On Three-Layer Data Markets

### v3

@@ 1 Introduction
With the advent of new functionalities by information technology companies such as Apple or Google that reduce or eliminate tracking via third-party cookies, there has been an increased reliance on first-party data, which is information gathered by a company from its own user base. For example, as elaborated by Cross [2023b], major retailers such as BestBuy and Walmart capitalize on this data through loyalty programs, customer purchase records, and subscriptions. They do this by selling the data to advertisers who aim to place targeted ads on the retailers’ websites or across various online platforms. In a similar vein, Cross [2023a] notes that Mastercard, a leading payment technology firm, trades cardholder transaction information via third-party online data marketplaces and its internal “Data and Services” division. This practice grants numerous entities access to extensive consumer data and insights. For instance, their Intelligent Targeting service allows businesses to leverage “Mastercard 360 data insights” to create and execute advertising campaigns that specifically target potential “high-value” customers.
@@
In this complex digital ecosystem, users face the dilemma of benefiting from the services provided by these platforms while potentially compromising their privacy. Platforms, on the other hand, must balance the profitability derived from data sharing with the need to maintain user trust and satisfaction. This interplay between users’ privacy concerns and the economic incentives of platforms raises important questions about the sustainability of current data market practices. Notable regulatory actions have been taken to improve user experiences. For instance, the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) have been introduced to protect user privacy and to limit the amount of data shared with the platforms. Although, at first sight, this helps users, there have been several studies pointing to its detrimental impact both on platforms and users. This is because these regulations have led to a more concentrated market structure by limiting competition in data markets. This results from the regulation’s restrictions on data sharing, which may slow down data-based innovations and services. For instance, studies have shown that GDPR has led to a significant reduction in the number of available apps in the market, with new app entries falling by half following the regulation’s implementation [see, e.g., Aridor et al., 2021, Janssen et al., 2022]. This decline in market entries indicates a potential loss of innovation as new and smaller players find it increasingly difficult to compete under these regulations. This effect is particularly detrimental to small companies that rely on these data synergies for innovation and market growth [Gal and Aviv, 2020]. On a similar note, Demirer et al. [2024] also show that GDPR, on average, represents a 20% increase in the cost of data, with the “distortionary effects of the GDPR are highest for the smallest firms.”
@@ 1.1 Our model and main results
We model the interaction between the players as a multi-stage game in which, in the first stage, the platforms decide whether they want to provide services to the user and the privacy level they want to deliver to the data buyer (so that the user is incentivized to share). We model the privacy level that the platform provides as a noise level added to the user data. Adding a suitable level of noise is the common practice for achieving privacy, primarily motivated by the differential privacy framework of Dwork et al. [2014]. This approach is widely used both in the machine learning literature and in real-world applications, ranging from census data [see, e.g., Abowd, 2018] to technology companies such as Google and Apple [see, e.g., Erlingsson et al., 2014, differential privacy team, 2017].
@@ 1.2 Related literature
Our paper broadly relates to the emerging literature on data markets and online platform behavior. Earlier works on this topic include Acquisti et al. [2016], Posner and Weyl [2018], Ali et al. [2019], Jones and Tonetti [2020], and Dosis and Sand-Zantman [2022], all of whom study the consequences of protecting and disclosing personal information about individuals. Additionally, Acemoglu et al. [2022], Bergemann et al. [2020], Ichihashi [2020b], and Fainmesser et al. [2022] study the privacy consequences of data externality (whereby a user’s data reveals information about others). Acemoglu et al. [2023a] study user-optimal privacy-preserving mechanisms that platforms can adopt to balance privacy and learning and Acemoglu et al. [2023b] develop an experimentation game to study the implications of the platform’s information advantage in product offering. Finally, Ichihashi [2021] studies data intermediaries, and Ichihashi [2020a] studies consumer privacy choices while online platforms dynamically adjust their privacy guarantees to incentivize more consumer data sharing.
@@
More closely related to ours are papers that study data sharing among users, platforms, and third-party data buyers and, in particular, the impacts of banning such data sharing. In this regard, Bimpikis et al. [2021] and Argenziano and Bonatti [2023] develop a two-round model of the interaction between a user and two competing platforms that can collect data on user preferences (and potentially use it for price discrimination). Bimpikis et al. [2021] show that banning data sharing can hurt the user when the platforms offer complementary products, and Argenziano and Bonatti [2023] show that banning data sharing can hurt the user if the benefits of knowing user preferences by the seller and personalized services are limited. The above papers consider the interactions between a user and a platform that itself uses the user data for price discrimination and, therefore, can potentially hurt the user. In contrast, we study a different problem that relates to how platforms sell user data to other third-party data buyers. Therefore, in addition to the differences in modeling choices and analysis, we depart from the above papers by considering a three-layer data market that includes the interaction between the platforms and the data buyer and focusing on deriving insights on effective regulations. Furthermore, Fang and Kim [2021] and Bergemann and Bonatti [2024] develop a model to analyze the impact of competition on how data-rich platforms can extract value from sellers through advertising.
@@
Also related to our work is Ravichandran and Korula [2019], who empirically study the effect of the data-sharing ban on platforms’ revenues, and Madsen and Vellodi [2023], who study the effects of a complete ban on innovation (we refer to Pino [2022] for a comprehensive survey on the microeconomics of data). We depart from these papers by focusing on the impact of regulations (and in particular, data-sharing ban) on users’ utility. Our paper also relates to the literature that studies various forms of monetizing user data by platforms, such as selling cookies [Bergemann and Bonatti, 2015], efficient pricing for large data sets [Abolhassani et al., 2017], dynamic sales [Immorlica et al., 2021], [Drakopoulos and Makhdoumi, 2023], monetization while ensuring the data is replicable [Falconer et al., 2023], and the design and price of information where a data buyer faces a decision problem under uncertainty and can augment their initial private information with supplemental data from a data seller [Bergemann et al., 2018].
@@
More broadly, our paper relates to the growing literature on data acquisition and the design of mechanisms for incentivizing data sharing [Ghosh and Roth, 2011, Ligett and Roth, 2012, Nissim et al., 2014, Cummings et al., 2015, Chen et al., 2018, Chen and Zheng, 2019, Fallah et al., 2023, Cummings et al., 2023, Fallah et al., 2022, Karimireddy et al., 2022]. We depart from this line of work as we consider a three-layer market where the privacy cost occurs due to selling data to a third-party buyer.
@@
Finally, our paper relates to the supply chain models with competition in the intermediate layer. In these models, upstream and downstream firms interact through contractual arrangements, and the middle layer competes either in prices or quantities (see, e.g., Section 5 of [Cachon, 2003]). Related work by Lariviere [1999], Corbett et al. [2004], and Bernstein and Federgruen [2005] explores how such competition shapes equilibrium prices, efficiency, and welfare. In these settings, competition among intermediaries tends to benefit the downstream buyer and can reduce the intermediaries’ profits—an observation that parallels our finding that platform competition benefits the data buyer but not the users. Despite this structural analogy, our framework departs from the supply chain literature in several key ways. First, the externality in our model arises from information leakage and privacy costs rather than inventory or production risk, leading to a fundamentally different welfare trade-off. Second, while classical models feature linear or convex demand and cost functions, the data buyer’s valuation in our model is submodular across platforms, capturing diminishing marginal returns to heterogeneous data sources. Finally, instead of price or quantity coordination, the platforms strategically choose their information quality (noise levels) and prices, mediating user participation and buyer demand.
@@ 2 Model
for the price $p_{i}$, where $\sigma_{i}^{2}$ represents the variance of the noise that platform $i$ adds to the data they share with the data buyer. In light of the discussion on differential privacy in the introduction, adding Gaussian noise—known as the Gaussian mechanism—makes the user’s data locally differentially private. This means that a change in a user’s data has a limited impact on the distribution of the shared signal $\tilde{s}_{i}$, where the extent of the impact is controlled by the variance $\sigma_{i}^{2}$. This, in turn, limits how much information the signal $\tilde{s}_{i}$ reveals about the user’s data to the data buyer. [^2] More formally, a randomized mapping $\mathcal{M}(\cdot)$ is said to be $(\varepsilon,\delta)$-locally differentially private if $\mathbb{P}(\mathcal{M}(s)\in W)\leq\exp(\varepsilon)\mathbb{P}(\mathcal{M}(s^{\prime})\in W)+\delta,$ for any $s,s^{\prime}$ and any measurable set $W$. See Duchi [2016] for further details and the characterization of the Gaussian mechanism.
@@ 2.1 Utility functions
Measuring the revealed information, or information gain, through the variance of the posterior distribution has been used widely in economics, statistics, and learning theory [see, e.g., Breiman et al., 1984, Bergemann et al., 2020, Acemoglu et al., 2022].
@@ 3 Preliminary Equilibrium Characterization
For a platform strategy profile $(\bm{\sigma},\bm{e})$, we select the platforms’ Pareto-optimal one if there are multiple user equilibria. Using Lemma 5, such a user equilibrium exists and is the one that shares with the highest number of platforms. [^7] We can view this as Stackelberg Nash equilibrium in which because the platforms move first, we break the ties in their favor. Similar concepts have appeared in the literature with different terms, such as sender-preferred equilibrium in Bayesian persuasion [Kamenica and Gentzkow, 2011].
@@ 5 Equilibrium Characterization with $K>2$ Platforms
We are interested in the second regime where $\alpha$ is large enough so that the user is not incentivized to share with the platform, irrespective of the noise level choice of the platform. In this case, as established in Proposition 6, both the user and the data buyer utilities are zero. Now, let us compare this setting to a setting with $K\geq 2$ platforms that are competing to obtain the user’s data and sell it to the data buyer, established in Theorem 1. Interestingly, we observe that competition does not help the user (as the user obtains zero utility for any $K$). However, increasing the number of platforms from $K=1$ to $K\geq 2$ improves the data buyer’s utility. To gain the intuition for this observation, notice that the user’s utility will always be zero because otherwise, the platforms can always decrease their noise levels by a small margin so that the user still shares and increases their payments. Now, why does the competition help the data buyer? This is because of the data externality among the platforms’ data about the user, similar to that of Acemoglu et al. [2022] and Bergemann et al. [2020]. In particular, the data of a platform (because of the submodularity of the revealed information, proved in Lemma 3) decreases the marginal value of another platform’s data, and therefore, the platforms end up competing with each other, and sell their data at lower marginal prices.
@@ 7 Conclusion
There are several exciting avenues for future research. In our framework, we have modeled a user as a representative of a market segment. Extending our analysis to a setting with multiple heterogeneous users would be very worthwhile. This is because data externalities among different users may cause indirect information disclosure, which in turn could alter both user strategies and potential regulations aimed at enhancing user experience. Similarly, our focus has been on a representative data buyer who aims to acquire user data and learn their preferences. It would be useful to consider an elaborated model with multiple buyers, given that user data is inherently multi-dimensional and can be aggregated from diverse sources, potentially modifying user strategies and calling for tailored regulation. Lastly, we have investigated regulation that restricts the transfer of user data to buyers. The challenge of developing alternative regulatory frameworks that address the intricacies of the data market—frameworks that are both adaptable and resilient enough to keep pace with swift technological advancements—will remain of significant importance for quite some timem to come.

### v4

@@ 1 Introduction
With the advent of new functionalities by information technology companies such as Apple or Google that reduce or eliminate tracking via third-party cookies, there has been an increased reliance on first-party data, which is information gathered by a company from its own user base. For example, as elaborated by cross2023new, major retailers such as BestBuy and Walmart capitalize on this data through loyalty programs, customer purchase records, and subscriptions. They do this by selling the data to advertisers who aim to place targeted ads on the retailers’ websites or across various online platforms. In a similar vein, Cross_2023 notes that Mastercard, a leading payment technology firm, trades cardholder transaction information via third-party online data marketplaces and its internal “Data and Services” division. This practice grants numerous entities access to extensive consumer data and insights. For instance, their Intelligent Targeting service allows businesses to leverage “Mastercard 360 data insights” to create and execute advertising campaigns that specifically target potential “high-value” customers.
@@
In this complex digital ecosystem, users face the dilemma of benefiting from the services provided by these platforms while potentially compromising their privacy. Platforms, on the other hand, must balance the profitability derived from data sharing with the need to maintain user trust and satisfaction. This interplay between users’ privacy concerns and the economic incentives of platforms raises important questions about the sustainability of current data market practices. Notable regulatory actions have been taken to improve user experiences. For instance, the General Data Protection Regulation (GDPR) and the California Consumer Privacy Act (CCPA) have been introduced to protect user privacy and to limit the amount of data shared with the platforms. Although, at first sight, this helps users, there have been several studies pointing to its detrimental impact both on platforms and users. This is because these regulations have led to a more concentrated market structure by limiting competition in data markets. This results from the regulation’s restrictions on data sharing, which may slow down data-based innovations and services. For instance, studies have shown that GDPR has led to a significant reduction in the number of available apps in the market, with new app entries falling by half following the regulation’s implementation [see, e.g., aridor2021effect, janssen2022gdpr]. This decline in market entries indicates a potential loss of innovation as new and smaller players find it increasingly difficult to compete under these regulations. This effect is particularly detrimental to small companies that rely on these data synergies for innovation and market growth [gal2020competitive]. On a similar note, demirer2024data also show that GDPR, on average, represents a 20% increase in the cost of data, with the “distortionary effects of the GDPR are highest for the smallest firms.”
@@ 1.1 Our model and main results
We model the interaction between the players as a multi-stage game in which, in the first stage, the platforms decide whether they want to provide services to the user and the privacy level they want to deliver to the data buyer (so that the user is incentivized to share). We model the privacy level that the platform provides as a noise level added to the user data. Adding a suitable level of noise is the common practice for achieving privacy, primarily motivated by the differential privacy framework of dwork2014algorithmic. This approach is widely used both in the machine learning literature and in real-world applications, ranging from census data [see, e.g., abowd2018us] to technology companies such as Google and Apple [see, e.g., erlingsson2014rappor, apple].
@@ 1.2 Related literature
Our paper broadly relates to the emerging literature on data markets and online platform behavior. Earlier works on this topic include acquisti2016economics, posner2018radical, ali2019voluntary, jones2020nonrivalry, and dosis2022, all of whom study the consequences of protecting and disclosing personal information about individuals. Additionally, acemoglu2019too, bergemann2020economics, ichihashi2020online, and fainmesser2022digital study the privacy consequences of data externality (whereby a user’s data reveals information about others). acemoglu2023good study user-optimal privacy-preserving mechanisms that platforms can adopt to balance privacy and learning and acemoglu2023model develop an experimentation game to study the implications of the platform’s information advantage in product offering. Finally, ichihashi2021competing studies data intermediaries, and ichihashi2020dynamic studies consumer privacy choices while online platforms dynamically adjust their privacy guarantees to incentivize more consumer data sharing.
@@
More closely related to ours are papers that study data sharing among users, platforms, and third-party data buyers and, in particular, the impacts of banning such data sharing. In this regard, bimpikis2021data and Argenziano2023 develop a two-round model of the interaction between a user and two competing platforms that can collect data on user preferences (and potentially use it for price discrimination). bimpikis2021data show that banning data sharing can hurt the user when the platforms offer complementary products, and Argenziano2023 show that banning data sharing can hurt the user if the benefits of knowing user preferences by the seller and personalized services are limited. The above papers consider the interactions between a user and a platform that itself uses the user data for price discrimination and, therefore, can potentially hurt the user. In contrast, we study a different problem that relates to how platforms sell user data to other third-party data buyers. Therefore, in addition to the differences in modeling choices and analysis, we depart from the above papers by considering a three-layer data market that includes the interaction between the platforms and the data buyer and focusing on deriving insights on effective regulations. Furthermore, fang2021data and bergemann2024data develop a model to analyze the impact of competition on how data-rich platforms can extract value from sellers through advertising.
@@
Also related to our work is ravichandran2019effect, who empirically study the effect of the data-sharing ban on platforms’ revenues, and madsen2023insider, who study the effects of a complete ban on innovation (we refer to pino2022microeconomics for a comprehensive survey on the microeconomics of data). We depart from these papers by focusing on the impact of regulations (and in particular, data-sharing ban) on users’ utility. Our paper also relates to the literature that studies various forms of monetizing user data by platforms, such as selling cookies [bergemann2015selling], efficient pricing for large data sets [abolhassani2017market], dynamic sales [immorlica2021buying], [drakopoulos2023providing], monetization while ensuring the data is replicable [falconer2023replicationrobust], and the design and price of information where a data buyer faces a decision problem under uncertainty and can augment their initial private information with supplemental data from a data seller [bergemann2018design].
@@
More broadly, our paper relates to the growing literature on data acquisition and the design of mechanisms for incentivizing data sharing [ghosh2011selling, ligett2012take, nissim2014redrawing, cummings2015accuracy, chen2018optimal, chen2019prior, fallah2023optimal, cummings2022optimal, fallah2022bridging, karimireddy2022mechanisms]. We depart from this line of work as we consider a three-layer market where the privacy cost occurs due to selling data to a third-party buyer.
@@
Finally, our paper relates to the supply chain models with competition in the intermediate layer. In these models, upstream and downstream firms interact through contractual arrangements, and the middle layer competes either in prices or quantities (see, e.g., Section 5 of [cachon2003supply]). Related work by lariviere1999supply, corbett2004designing, and bernstein2005decentralized explores how such competition shapes equilibrium prices, efficiency, and welfare. In these settings, competition among intermediaries tends to benefit the downstream buyer and can reduce the intermediaries’ profits—an observation that parallels our finding that platform competition benefits the data buyer but not the users. Despite this structural analogy, our framework departs from the supply chain literature in several key ways. First, the externality in our model arises from information leakage and privacy costs rather than inventory or production risk, leading to a fundamentally different welfare trade-off. Second, while classical models feature linear or convex demand and cost functions, the data buyer’s valuation in our model is submodular across platforms, capturing diminishing marginal returns to heterogeneous data sources. Finally, instead of price or quantity coordination, the platforms strategically choose their information quality (noise levels) and prices, mediating user participation and buyer demand.
@@ 2 Model
for the price $p_{i}$, where $\sigma_{i}^{2}$ represents the variance of the noise that platform $i$ adds to the data they share with the data buyer. In light of the discussion on differential privacy in the introduction, adding Gaussian noise—known as the Gaussian mechanism—makes the user’s data locally differentially private. This means that a change in a user’s data has a limited impact on the distribution of the shared signal $\tilde{s}_{i}$, where the extent of the impact is controlled by the variance $\sigma_{i}^{2}$. This, in turn, limits how much information the signal $\tilde{s}_{i}$ reveals about the user’s data to the data buyer. [^2] More formally, a randomized mapping $\mathcal{M}(\cdot)$ is said to be $(\varepsilon,\delta)$-locally differentially private if $\mathbb{P}(\mathcal{M}(s)\in W)\leq\exp(\varepsilon)\mathbb{P}(\mathcal{M}(s^{\prime})\in W)+\delta,$ for any $s,s^{\prime}$ and any measurable set $W$. See duchi2016lecture for further details and the characterization of the Gaussian mechanism.
@@ 2.1 Utility functions
Measuring the revealed information, or information gain, through the variance of the posterior distribution has been used widely in economics, statistics, and learning theory [see, e.g., breiman1984classification, bergemann2020economics, acemoglu2019too].
@@ 3 Preliminary Equilibrium Characterization
For a platform strategy profile $(\bm{\sigma},\bm{e})$, we select the platforms’ Pareto-optimal one if there are multiple user equilibria. Using Lemma 5, such a user equilibrium exists and is the one that shares with the highest number of platforms. [^7] We can view this as Stackelberg Nash equilibrium in which because the platforms move first, we break the ties in their favor. Similar concepts have appeared in the literature with different terms, such as sender-preferred equilibrium in Bayesian persuasion [kamenica2011bayesian].
@@ 5 Equilibrium Characterization with $K>2$ Platforms
We are interested in the second regime where $\alpha$ is large enough so that the user is not incentivized to share with the platform, irrespective of the noise level choice of the platform. In this case, as established in Proposition 6, both the user and the data buyer utilities are zero. Now, let us compare this setting to a setting with $K\geq 2$ platforms that are competing to obtain the user’s data and sell it to the data buyer, established in Theorem 1. Interestingly, we observe that competition does not help the user (as the user obtains zero utility for any $K$). However, increasing the number of platforms from $K=1$ to $K\geq 2$ improves the data buyer’s utility. To gain the intuition for this observation, notice that the user’s utility will always be zero because otherwise, the platforms can always decrease their noise levels by a small margin so that the user still shares and increases their payments. Now, why does the competition help the data buyer? This is because of the data externality among the platforms’ data about the user, similar to that of acemoglu2019too and bergemann2020economics. In particular, the data of a platform (because of the submodularity of the revealed information, proved in Lemma 3) decreases the marginal value of another platform’s data, and therefore, the platforms end up competing with each other, and sell their data at lower marginal prices.
@@ 7 Conclusion
There are several exciting avenues for future research. In our framework, we have modeled a user as a representative of a market segment. Extending our analysis to a setting with multiple heterogeneous users would be very worthwhile. This is because data externalities among different users may cause indirect information disclosure, which in turn could alter both user strategies and potential regulations aimed at enhancing user experience. Similarly, our focus has been on a representative data buyer who aims to acquire user data and learn their preferences. It would be useful to consider an elaborated model with multiple buyers, given that user data is inherently multi-dimensional and can be aggregated from diverse sources, potentially modifying user strategies and calling for tailored regulation. Lastly, we have investigated regulation that restricts the transfer of user data to buyers. The challenge of developing alternative regulatory frameworks that address the intricacies of the data market—frameworks that are both adaptable and resilient enough to keep pace with swift technological advancements—will remain of significant importance for quite some time to come.
8 Acknowledgment
Alireza Fallah acknowledges support from the European Research Council Synergy Program, the National Science Foundation under grant number DMS-1928930, and the Alfred P. Sloan Foundation under grant G-2021-16778. The latter two grants correspond to his residency at the Simons Laufer Mathematical Sciences Institute (formerly known as MSRI) in Berkeley, California, during the Fall 2023 semester. Michael Jordan acknowledges support from the Mathematical Data Science program of the Office of Naval Research under grant number N00014-21-1-2840 and the European Research Council Synergy Program.